OpenStudy (anonymous):

domain of (sqrt (t-62))^(1/3)

OpenStudy (anonymous):

all nonnegative real numbers. How do i write it in interval notation?

OpenStudy (amriju):

(\[[62,\infty]\]

OpenStudy (anonymous):

all real no such that x>=62

Parth (parthkohli):

@amriju Correction:\[\rm [62,+\infty)\]

Parth (parthkohli):

or,\[\rm \{x \in \mathbb R|x>61\}\]

OpenStudy (anonymous):

It was (-infinity, infinity) U [62, infinity) SIGH. Thank you though!